OpenStudy (dtan5457):

What is cot(theta)+tan(theta)=? csc2(theta) 2csc2(theta) sec2(theta) 2sec2(theta)

OpenStudy (mathstudent55):

\(\cot \theta + \tan \theta\) \(= \dfrac{\cos \theta}{\sin \theta} + \dfrac{\sin \theta}{\cos \theta} \) \(= \dfrac{\cos \theta}{\sin \theta} \times \dfrac{\cos \theta}{\cos \theta}+ \dfrac{\sin \theta}{\cos \theta} \times \dfrac{\sin \theta}{\sin \theta}\) \(= \dfrac{\cos^2 \theta}{\sin \theta \cos \theta} + \dfrac{\sin^2 \theta}{\sin \theta \cos \theta} \) \(= \dfrac{\cos^2 \theta + \sin^2 \theta}{\sin \theta \cos \theta}\) \(= \dfrac{1}{\sin \theta \cos \theta}\) \(= \sec \theta \csc \theta\) \(= 2 \csc 2 \theta\)